In some regions of west Africa, the frequency of the HBBS allele is 0.1. If this frequency is the result of a dynamic equilibrium due to the superior fitness of HBBSHBBA heterozygotes, and if HBBSHBBS homozygotes are essentially lethal, what is the intensity of selection against the HBBAHBBA homozygotes?

A) 0.91

B) 0.42

C) 0.33

D) 0.22

E) 0.11

Working solution would be appreciated
Answer #1

Cosidering heterozygous fitness or overdominance:

p= t/ (s+t)

Where:

p= allele frequency of A1 =0.1

s= selection coefficient of A1A1

t= selection coefficient of A2A2

For lethal allele, s=1

thus,

0.1 = (t)/ (1+t) = 0.11

Answer: E